# QG from SymQRG: AdS3/CFT2 Correspondence as Topological Symmetry-Preserving Quantum RG Flow

### v1

@@ Abstract
By analyzing the non-perturbative RG flows that explicitly preserve given symmetries, we demonstrate that they can be expressed as quantum path integrals of the SymTFT in one higher dimension. When the symmetries involved include Virasoro defect lines, such as in the case of $T\bar{T}$ deformations, the RG flow corresponds to the 3D quantum gravitational path integral. For each 2D CFT, we identify a corresponding ground state of the SymTFT, from which the Wheeler-DeWitt equation naturally emerges as a non-perturbative constraint. These observations are summarized in the slogan: SymQRG = QG. The recently proposed exact discrete formulation of Liouville theory in [1] allows us to identify a universal SymQRG kernel, constructed from quantum $6j$ symbols associated with $U_{q}(SL(2,\mathbb{R}))$, which manifests itself as an exact and analytical 3D background-independent MERA-type holographic tensor network. Many aspects of the AdS/CFT correspondence, including the factorization puzzle, admit a natural interpretation within this framework. This provides the first evidence suggesting that there is a universal holographic principle encompassing AdS/CFT and topological holography. We propose that the non-perturbative AdS/CFT correspondence is a maximal form of topological holography.
@@ 1 Introduction
A very important aspect of the AdS/CFT correspondence is that global symmetries of the field theory exhibit themselves as gauge symmetries in the bulk. In particular, conformal transformations of the field theory are connected to a subset of diffeomorphisms in the asymptotically AdSD+1 spacetime. Other global symmetries in the field theory also manifest as gauge symmetries in the holographic dual. This structure bears striking resemblance to the “topological holographic principle”[5, 6, 7, *Kong:2020cie, *Kong:2020jne, *Chatterjee:2022kxb, 11, 12, 13, 14, 15, 16, 17], proposed in the study of generalized symmetries[18], [^1] The modern approach to symmetries involves studying topological operators within the theory; hence, we will also refer to this as topological symmetries. where the path integral of a quantum field theory (QFT) with symmetries on a $D$-manifold $\Sigma_{D}$ is produced from the path integral of a topological field theory (often called SymTFT) in $(D+1)$ dimensions on a manifold of the form $\Sigma_{D}\times I$, with $I$ being an interval. This is often called “sandwich construction”[5, 7, 6, 11]. The SymTFT is chosen so that a given global symmetry of the QFT is gauged in the TQFT [19, 7]. This strongly suggests that the two holographic principles are closely related.
@@
Can the combination of topological symmetries and RG, the two most important non-perturbative aspects of strongly interacting QFTs, reveal new insights into quantum gravity? In this paper, we demonstrate that a “topological symmetry-preserving quantum RG” process, referred to as “SymQRG”, can be translated to the path integral of the SymTFT, which probes deeper into the bulk of asymptotically AdSD+1 spacetimes. In particular in 2D, RG processes preserving an irrational collection of topological lines related to Virasoro representations produce a precise 3D quantum gravitational path integral that is given by two copies of the Virasoro/Teichmuller TQFT[23, 24, 25, 1]. i.e. When the symmetry concerned contains an irrational collection of topological line operators[18, 23] related to Virasoro representations, the SymTFT is a long-range entangled topological order[26, 27, 28, 29, *PhysRevLett.96.110405] that contains 3D quantum gravity.
@@ 1.1 AdS/CFT correspondence from generalized symmetries
Translated to the AdS/CFT correspondence, on the CFT side, the ket state -or the topological boundary condition $\mathcal{B}$ together with the bulk-encodes all the sewing rules or operator product expansion (OPE) data of the 2D CFT[36, *Fuchs:2003id, *Fuchs:2003id1, *Fuchs:2004xi, *Fjelstad:2005ua, *Frohlich:2006ch, *Fjelstad:2006aw]. This state is inherently 3D. When the CFT contains a large number of degrees of freedom, the bulk thus emerging has semi-classical geometrical meanings. In the AdS context, instead of treating this as an explicit boundary condition, we can view it as a complete specification of the quantum measure for the bulk path integral, which is often obscured in the semi-classical limit. This precise choice of measure guarantees the exact correspondence between 2D CFT and 3D gravity, and ensures factorization in the bulk. In general, the bulk theory coming from a given CFT does not sum over all geometries. One of the main observations of this paper is the importance of an underlying 3D quantum state that underlies any 2D CFT, and it determines the geometries to be summed in the bulk. On the other hand, the bra state—or the physical boundary condition—defines theories at different scales along the symmetry-preserving RG trajectory. These states can also be interpreted as basis states at different finite cutoffs in the bulk. Together, the ket and bra states form the basis of SymQRG and bridge the AdS/CFT correspondence (QG).
@@
In the examples we study explicitly in this paper, ranging from rational CFT and the Liouville theory, the corresponding holographic duals preserving the symmetries are topological. However, in more general cases where the symmetries involve an infinite and perhaps even continuous spectrum of charges, the bulk “SymTFT” may no longer remain topological. See for example [47]. [^3] As we will see, our construction of the bulk and the topological boundary comes from the OPE coefficients of the theory, which captures all the dynamical information when we use the whole chiral algebra of the CFT. Thus, we think it is unlikely that it only captures a topological subsector of the theory. Perhaps SymTFT is more generally “SymFT”.
@@ 1.3 Summary of framework
    Given a 2D CFT, we first have a chiral algebra $V$(we mainly focus on Virasoro in this paper). The representations of $V$ form the (irrational generalization of the) Moore-Seiberg category $C=\text{Rep}(V)$ [52, *Moore:1989vd].
@@
    Using $C$, we can construct the 3D SymTFT of this theory using the Turaev-Viro/Levin-Wen string-net type construction[26, 27].
@@
    The bulk SymTFT is related to 3D gravity in either an exact or a coarse-grained sense, when the SymTFT contains a large number of Virasoro defect lines, corresponding to heavy operators in the CFT. Summing over certain quantum geometries in the bulk emerges from the algebraic data of the CFT.
@@ 2.2 Quantum RG
Since the discovery of the AdS/CFT correspondence, it has been proposed that the change of scale, or RG flow, in the $D$-dimensional boundary theory is closely tied to the emergence of the extra radial direction in $D+1$-dimensional asymptotically AdS spacetime [2, 3, 4, 20]. This connection has been studied systematically under the framework of “holographic RG flow”[21, 22]. However, as noted in [57, *Faulkner:2010jy, 59] and emphasized in [48], there is a notable distinction between holographic RG and conventional Wilson-Kadanoff RG. In traditional RG procedures, the couplings in the boundary theory evolve according to first-order differential equations dictated by the beta functions, and they remain classical/non-dynamical. In contrast, in holography, these couplings correspond to fields in the gravitational bulk, where they are inherently quantum/dynamical. Most importantly, in the AdS/CFT correspondence, the boundary metric tensor $h_{ij}$ serves as a classical background source for the stress energy tensor $T_{ij}$ of the boundary CFT. This fixed metric defines the geometric 2D background on which the CFT is placed, and the CFT stress tensor encodes the response of the theory to variations in this background. In contrast, bulk theory is a theory of quantum gravity, where the bulk metric tensor $g_{\mu\nu}$ is a dynamical field that evolves according to gravitational dynamics and is capable of quantum fluctuations.
To resolve this tension, Lee proposed a manipulation in [48, 49] that converts the sources into dynamical and quantum degrees of freedom in the bulk. This is achieved by retaining only a subset of operators that forms a basis of simple operators in the operator space during RG flows. Specifically, in large $N$ holographic theories, double-trace operators are inevitably generated from single-trace operators during RG transformations [57, *Faulkner:2010jy], and we re-express these double-trace operators in terms of single-trace operators and keep only the single-traces at each step of the RG. We will adopt this perspective in the context of the $T\bar{T}$ deformation in Sec. 3, explicitly demonstrating how this procedure transforms the boundary metric $h_{ij}$ into the dynamical and quantum metric fields $g_{ij}$ in the bulk. This approach, known as Quantum RG, has been shown to reproduce key aspects of the quantum holographic correspondence in specific settings.
@@ 2.3 SymQRG: Topological Symmetry-Preserving Quantum RG
In this paper, we take a step further to explain the meaning of this projection into simple operators from a symmetry principle that extends beyond the large $N$ limit, highlighting its crucial role in the context of the AdS3/CFT2 correspondence. The key insight is that holography can be understood as an RG transformation that explicitly preserves certain topological symmetries of the 2D theories. This collection of symmetries includes generically non-invertible symmetries, and they follow from the representation of the chiral algebra inherent in any 2D CFTs, as first discovered by Moore and Seiberg[60, 52, *Moore:1989vd]. These topological symmetries manifest themselves as topological line operators in the CFT[23, 61, 62]. The connection between this symmetry and the complete set of CFT gluing rules was elucidated in a series of seminal works collectively referred to the FRS construction in the literature, named after Fuchs, Runkel and Schweigert, as well as Fröhlich and Fjelstad [36, *Fuchs:2003id, *Fuchs:2003id1, *Fuchs:2004xi, *Fjelstad:2005ua, *Frohlich:2006ch, *Fjelstad:2006aw]. The emergence of a 3D hyperbolic geometry from 2D CFTs is precisely captured by this “Virasoro modular geometry” for irrational CFTs[43, 63, 64, 65, 66]. More details on the emergence of the quantum group/non-commutative geometry $U_{q}(SL(2,\mathbb{R}))$ and its connection to 3D gravity are the focus of Sec. 5.
@@
where the subscript $\Lambda$ denotes some reference scale (which we will make precise in explicit constructions), $\gamma_{ij}$ is the metric on which the theory is defined and $\lambda$ labels the dimensionful couplings of the theory. In Sec. 3, $\lambda$ refers to the coupling for the $T\bar{T}$ deformation, which has mass dimension $-2$. This is in turn related to the size of the holes we introduce on the lattice in Sec. 4. In the language of generalized symmetries[18], this overlap realizes the “sandwich construction” [5, 7, 6, 11], as shown in Fig. 4, where ${}_{\Lambda}\bra{\Omega^{\lambda}(\gamma_{ij})}$ | corresponds to a physical boundary condition that captures geometrical notions such as distances and positions in a 2D QFT. For instance, in our $T\bar{T}$ example in Sec. 3.3, this state can be interpreted as the finite cutoff “position space” basis states [68, *PhysRevD.28.2960] in 3D quantum gravity; In Sec. 4, we use the Hilbert space of the Levin-Wen string-net model[27] underlying the tensor network construction to write down a continuous family of states, labelled by the size of the holes. In both of these cases, $\lambda=0$ corresponds to an RG fixed point (CFT). We also want to emphasize that these states describes the kinematical aspects of the 2D theory. Dynamics are encoded in the ket state |$\ket{\Psi}_{\Lambda}$, which corresponds to the state obtained from a 3D path integral associated with a topological boundary condition [^6] This combination is referred to as a “quiche” in [6].. The state contains the algebraic information of 2D CFT, such as its sewing rules, i.e. all the OPE coefficients. [36, *Fuchs:2003id, *Fuchs:2003id1, *Fuchs:2004xi, *Fjelstad:2005ua, *Fjelstad:2006aw]. The holographic principle thus equates this 3D wavefunction with 2D partition functions, and we will elaborate more on this in Sec. 6.
@@
where we keep the time dependence explicit to make the relation to (2.12) manifest. The third equality returns to the basis defined at $t=t_{0}$, which also coincides with the basis state in the Schrödinger picture, using the Hamiltonian $\hat{H}$. The kernel $K(x,x^{\prime},t,t_{0})$ in the fourth line is the transition amplitude, which can be expressed explicitly using Feynman’s path integral. We will see in Sec. 3.1, the quantum RG procedure[48, 49] is exactly analogous to time evolution in quantum mechanics. The position $x$ is the fixed argument of our wavefunction $\Psi(x)$, similar to the role played by the source $J$ or $h$ in $Z(J)$ or $Z(h)$ in QFTs. In fact, solution to the $T\bar{T}$ deformation exactly takes the form of (2.12), and this deformed partition function is nothing but the Wheeler-DeWitt wave functional in the Heisenberg picture. This comparison with usual quantum mechanical treatment makes manifest the existence of an underlying quantum state |$\ket{\Psi}_{\Lambda}$, which we will show to be a long-range entangled ground state of the topological order in the bulk, associated to the underlying Moore-Seiberg data. The ${}_{\Lambda}\bra{\Omega^{\lambda}(\gamma_{ij})}$ | states correspond to finite cutoff Dirichlet boundary conditions in the AdS/CFT correspondence. Alternatively, these states can be understood as the implementation of mixed boundary conditions at asymptotic infinity [^9] Naively, the bulk is a TQFT, so the notion of radial distances in the bulk does not seem to make sense. However, this becomes possible because the boundary conditions break topological invariance, and there is a scale $\lambda$ within these boundary conditions. The relative scale between the parameters of different boundary conditions can then be interpreted as the parameterization of a radial distance, arising from double-trace deformations[70, 71, 72, 73, *Guica:2022gts, 75, *Ebert:2022ehb]. Interestingly, when interpreted as boundary conditions, this suggests that these boundary conditions all generate the same Virasoro algebra in a nontrivial way[73, *Guica:2022gts].

### v2

@@ Abstract
By analyzing the non-perturbative RG flows that explicitly preserve given symmetries, we demonstrate that they can be expressed as quantum path integrals of the SymTFT in one higher dimension. When the symmetries involved include Virasoro defect lines, such as in the case of $T\bar{T}$ deformations, the RG flow corresponds to the 3D quantum gravitational path integral. For each 2D CFT, we identify a corresponding ground state of the SymTFT, from which the Wheeler-DeWitt equation naturally emerges as a non-perturbative constraint. These observations are summarized in the slogan: SymQRG = QG. The recently proposed exact discrete formulation of Liouville theory in [1] allows us to identify a universal SymQRG kernel, constructed from quantum $6j$ symbols associated with $U_{q}(SL(2,\mathbb{R}))$. This kernel is directly related to the quantum path integral of the Virasoro TQFT, and manifests as an exact and analytical 3D background-independent MERA-type holographic tensor network. Many aspects of the AdS/CFT correspondence, including the factorization puzzle, admit a natural interpretation within this framework. This provides the first evidence suggesting that there is a universal holographic principle encompassing AdS/CFT and topological holography. We propose that the non-perturbative AdS/CFT correspondence is a maximal form of topological holography.
@@ 1 Introduction
A very important aspect of the AdS/CFT correspondence is that global symmetries of the field theory exhibit themselves as gauge symmetries in the bulk. In particular, conformal transformations of the field theory are connected to a subset of diffeomorphisms in the asymptotically AdSD+1 spacetime. Other global symmetries in the field theory also manifest as gauge symmetries in the holographic dual. This structure bears striking resemblance to the “topological holographic principle” , proposed in the study of generalized symmetries[5], [^1] The modern approach to symmetries involves studying topological operators within the theory; hence, we will also refer to this as topological symmetries. where the path integral of a quantum field theory (QFT) with symmetries on a $D$-manifold $\Sigma_{D}$ is produced from the path integral of a topological field theory (often called SymTFT) in $(D+1)$ dimensions on a manifold of the form $\Sigma_{D}\times I$, with $I$ being an interval. This is often called “sandwich construction”[6, 7, 8, 9, 10]. The SymTFT is chosen so that a given global symmetry of the QFT is gauged in the TQFT [11, 7]. This strongly suggests that the two holographic principles are closely related.
@@
Can the combination of topological symmetries and RG, the two most important non-perturbative aspects of strongly interacting QFTs, reveal new insights into quantum gravity? In this paper, we demonstrate that a “topological symmetry-preserving quantum RG” process, referred to as “SymQRG”, can be translated to the path integral of the SymTFT, which probes deeper into the bulk of asymptotically AdSD+1 spacetimes. In particular in 2D, RG processes preserving an irrational collection of topological lines related to Virasoro representations produce a precise 3D quantum gravitational path integral that is given by two copies of the Virasoro/Teichmuller TQFT[15, 16, 17, 1]. i.e. When the symmetry concerned contains an irrational collection of topological line operators[5, 15] related to Virasoro representations, the SymTFT is a long-range entangled topological order[18, 19, 20, 21, *PhysRevLett.96.110405] that contains 3D quantum gravity.
@@ 1.1 AdS/CFT correspondence from generalized symmetries
Translated to the AdS/CFT correspondence, on the CFT side, the ket state -or the topological boundary condition $\mathcal{B}$ together with the bulk-encodes all the sewing rules or operator product expansion (OPE) data of the 2D CFT[28, *Fuchs:2003id, *Fuchs:2003id1, *Fuchs:2004xi, *Fjelstad:2005ua, *Frohlich:2006ch, *Fjelstad:2006aw]. This state is inherently 3D. When the CFT contains a large number of degrees of freedom, the bulk thus emerging has semi-classical geometrical meanings. In the AdS context, instead of treating this as an explicit boundary condition, we can view it as a complete specification of the quantum measure for the bulk path integral, which is often obscured in the semi-classical limit. This precise choice of measure guarantees the exact correspondence between 2D CFT and 3D gravity, and ensures factorization in the bulk. In general, the bulk theory coming from a given CFT does not sum over all geometries. One of the main observations of this paper is the importance of an underlying 3D quantum state that underlies any 2D CFT, and it determines the geometries to be summed in the bulk. On the other hand, the bra state—or the physical boundary condition—defines theories at different scales along the symmetry-preserving RG trajectory. These states can also be interpreted as basis states at different finite cutoffs in the bulk. Together, the ket and bra states form the basis of SymQRG and bridge the AdS/CFT correspondence (QG).
@@
In the examples we study explicitly in this paper, ranging from rational CFT and the Liouville theory, the corresponding holographic duals preserving the symmetries are topological. However, in more general cases where the symmetries involve an infinite and perhaps even continuous spectrum of charges, the bulk “SymTFT” may no longer remain topological. See for example [39]. [^3] As we will see, our construction of the bulk and the topological boundary comes from the OPE coefficients of the theory, which captures all the dynamical information when we use the chiral algebra (Virasoro algebra) of the CFT. Thus, we think it is unlikely that it only captures a topological subsector of the theory. Perhaps SymTFT is more generally “SymFT”.
@@ 1.3 Summary of framework
    Given a non-chiral 2D CFT, we first have a chiral algebra $V$. The representations of $V$ form the (irrational generalization of the) Moore-Seiberg category $C=\text{Rep}(V)$ [44, *Moore:1989vd].
@@
    Using $C$, we can construct the 3D SymTFT of this theory, for example using the Turaev-Viro/Levin-Wen string-net type construction[18, 19].
@@
    The 3D bulk theory that captures the maximal amount of symmetries of the boundary theory encodes all topological defects that commute only with the Virasoro algebra of the 2D CFT. [^4] Topological defect lines that commute with extended chiral algebras form a subset of those that commute only with the Virasoro algebra. It becomes related to 3D gravity, either in an exact or coarse-grained sense, when the SymTFT contains a large number of Virasoro defect lines, corresponding to heavy operators in the CFT. The summation over certain quantum geometries in the bulk naturally emerges from the algebraic data of the CFT.
@@ 2.2 Quantum RG
Since the discovery of the AdS/CFT correspondence, it has been proposed that the change of scale, or RG flow, in the $D$-dimensional boundary theory is closely tied to the emergence of the extra radial direction in $D+1$-dimensional asymptotically AdS spacetime [2, 3, 4, 12]. This connection has been studied systematically under the framework of “holographic RG flow”[13, 14]. However, as noted in [49, *Faulkner:2010jy, 51] and emphasized in [40], there is a notable distinction between holographic RG and conventional Wilson-Kadanoff RG. In traditional RG procedures, the couplings in the boundary theory evolve according to first-order differential equations dictated by the beta functions, and they remain classical/non-dynamical. In contrast, in holography, these couplings correspond to fields in the gravitational bulk, where they are inherently quantum/dynamical. Most importantly, in the AdS/CFT correspondence, the boundary metric tensor $h_{ij}$ serves as a classical background source for the stress energy tensor $T_{ij}$ of the boundary CFT. This fixed metric defines the geometric 2D background on which the CFT is placed, and the CFT stress tensor encodes the response of the theory to variations in this background. In contrast, bulk theory is a theory of quantum gravity, where the bulk metric tensor $g_{\mu\nu}$ is a dynamical field that evolves according to gravitational dynamics and is capable of quantum fluctuations.
To resolve this tension, Lee proposed a manipulation in [40, 41] that converts the sources into dynamical and quantum degrees of freedom in the bulk. This is achieved by retaining only a subset of operators that forms a basis of simple operators in the operator space during RG flows. Specifically, in large $N$ holographic theories, double-trace operators are inevitably generated from single-trace operators during RG transformations [49, *Faulkner:2010jy], and we re-express these double-trace operators in terms of single-trace operators and keep only the single-traces at each step of the RG. We will adopt this perspective in the context of the $T\bar{T}$ deformation in Sec. 3, explicitly demonstrating how this procedure transforms the boundary metric $h_{ij}$ into the dynamical and quantum metric fields $g_{ij}$ in the bulk. This approach, known as Quantum RG, has been shown to reproduce key aspects of the quantum holographic correspondence in specific settings.
@@ 2.3 SymQRG: Topological Symmetry-Preserving Quantum RG
2.3.1 General framework
In this paper, we take a step further to explain the meaning of this projection into simple operators from a symmetry principle that extends beyond the large $N$ limit, highlighting its crucial role in the context of the AdS3/CFT2 correspondence. The key insight is that holography can be understood as an RG transformation that explicitly preserves certain topological symmetries of the 2D theories. This collection of symmetries includes generically non-invertible symmetries, and they follow from the (irrational generalization of) representations of the chiral algebra inherent in any 2D CFTs, as first discovered by Moore and Seiberg[52, 44, *Moore:1989vd]. These topological symmetries manifest themselves as topological line operators in the CFT[15, 53, 54, 55, 56]. The connection between this symmetry and the complete set of CFT gluing rules was elucidated in a series of seminal works collectively referred to the FRS construction in the literature, named after Fuchs, Runkel and Schweigert, as well as Fröhlich and Fjelstad [28, *Fuchs:2003id, *Fuchs:2003id1, *Fuchs:2004xi, *Fjelstad:2005ua, *Frohlich:2006ch, *Fjelstad:2006aw]. The emergence of a 3D hyperbolic geometry from 2D CFTs is precisely captured by this “Virasoro modular geometry” for irrational CFTs[35, 57, 58, 59, 60]. More details on the emergence of the quantum group/non-commutative geometry $U_{q}(SL(2,\mathbb{R}))$ and its connection to 3D gravity are the focus of Sec. 5.
@@ 2.3.1 General framework
where the subscript $\Lambda$ denotes some reference scale (which we will make precise in explicit constructions), $\gamma_{ij}$ is the metric on which the theory is defined and $\lambda$ labels the dimensionful couplings of the theory. In Sec. 3, $\lambda$ refers to the coupling for the $T\bar{T}$ deformation, which has mass dimension $-2$. This is in turn related to the size of the holes we introduce on the lattice in Sec. 4. In the language of generalized symmetries[5], this overlap realizes the “sandwich construction” [6, 7, 8, 9, 10], as shown in Fig. 4, where ${}_{\Lambda}\bra{\Omega^{\lambda}(\gamma_{ij})}$ | corresponds to a physical boundary condition that captures geometrical notions such as distances and positions in a 2D QFT. For instance, in our $T\bar{T}$ example in Sec. 3.3, this state can be interpreted as the finite cutoff “position space” basis states [62, *PhysRevD.28.2960] in 3D quantum gravity; In Sec. 4, we use the Hilbert space of the Levin-Wen string-net model[19] underlying the tensor network construction to write down a continuous family of states, labelled by the size of the holes. In both of these cases, $\lambda=0$ corresponds to an RG fixed point (CFT). We also want to emphasize that these states describes the kinematical aspects of the 2D theory. Dynamics are encoded in the ket state |$\ket{\Psi}_{\Lambda}$, which corresponds to the state obtained from a 3D path integral associated with a topological boundary condition [^7] This combination is referred to as a “quiche” in [8].. The state contains the algebraic information of 2D CFT, such as its sewing rules, i.e. all the OPE coefficients. [28, *Fuchs:2003id, *Fuchs:2003id1, *Fuchs:2004xi, *Fjelstad:2005ua, *Fjelstad:2006aw]. The holographic principle thus equates this 3D wavefunction with 2D partition functions, and we will elaborate more on this in Sec. 6.
@@
where we keep the time dependence explicit to make the relation to (2.12) manifest. The third equality returns to the basis defined at $t=t_{0}$, which also coincides with the basis state in the Schrödinger picture, using the Hamiltonian $\hat{H}$. The kernel $K(x,x^{\prime},t,t_{0})$ in the fourth line is the transition amplitude, which can be expressed explicitly using Feynman’s path integral. We will see in Sec. 3.1, the quantum RG procedure[40, 41] is exactly analogous to time evolution in quantum mechanics. The position $x$ is the fixed argument of our wavefunction $\Psi(x)$, similar to the role played by the source $J$ or $h$ in $Z(J)$ or $Z(h)$ in QFTs. In fact, solution to the $T\bar{T}$ deformation exactly takes the form of (2.12), and this deformed partition function is nothing but the Wheeler-DeWitt wave functional in the Heisenberg picture. This comparison with usual quantum mechanical treatment makes manifest the existence of an underlying quantum state |$\ket{\Psi}_{\Lambda}$, which we will show to be a long-range entangled ground state of the topological order in the bulk, associated to the underlying Moore-Seiberg data. The ${}_{\Lambda}\bra{\Omega^{\lambda}(\gamma_{ij})}$ | states correspond to finite cutoff Dirichlet boundary conditions in the AdS/CFT correspondence. Alternatively, these states can be understood as the implementation of mixed boundary conditions at asymptotic infinity [^10] Naively, the bulk is a TQFT, so the notion of radial distances in the bulk does not seem to make sense. However, this becomes possible because the boundary conditions break topological invariance, and there is a scale $\lambda$ within these boundary conditions. The relative scale between the parameters of different boundary conditions can then be interpreted as the parametrization of a radial distance, arising from double-trace deformations[64, 65, 66, 67, *Guica:2022gts, 69, *Ebert:2022ehb]. Interestingly, when interpreted as boundary conditions, this suggests that these boundary conditions all generate the same Virasoro algebra in a nontrivial way[67, *Guica:2022gts].
@@
2.3.2 Illustrative Example
To illustrate the idea, we consider a simple example related to the torus partition functions of 2D rational CFTs. On a torus, the CFT partition function is given by
$Z(\tau,\bar{\tau})=\sum_{I,J}Z_{IJ}\chi_{I}(\tau)\bar{\chi}_{J}(\bar{\tau})~{},$ (2.14)
where $(I,J)$ label conformal primaries in the CFT spectrum, with chiral/anti-chiral representations $I$ and $J$. $\chi_{I}(\tau)$ is the associated character with modulus $\tau$. As shown in [72, 73], the Hilbert space of torus conformal blocks is isomorphic to that of a 3D TQFT on a torus, leading to a 3D quantum state
|$\ket{\Psi}=\sum_{I,J}Z_{IJ}\ket{I,J}~{},$ (2.15)
where |$\ket{I,J}$ can be constructed from 3D solid torus path integral with Wilson line insertions.
This quantum state encodes the operators appearing in the spectrum of the 2D CFT. It can be constructed through the insertion of a mesh of Wilson lines and alternatively described by a topological boundary condition of the 3D TQFT, as we will explain in Sec. 6.
The coefficients $Z_{IJ}$ represent algebraic/topological data of the CFT, ensuring a modular invariant spectrum. On more general manifolds, they are explicitly related to closed CFT OPE coefficients via the closed pair-of-pants decomposition. The observation in [28] is that these coefficients can also be constructed from BCFT open OPE coefficients, a fact we will leverage to build the exact CFT tensor network. Notably, these are numbers independent of the moduli and are intrinsically topological in nature.
To recover the CFT partition function with moduli dependence, we can construct another physical boundary condition in the 3D TQFT, corresponding to a quantum state labeled by $\bra{\tau,\bar{\tau}}$ |, such that
$\bra{\tau,\bar{\tau}}I,J\rangle=\chi_{I}(\tau)\bar{\chi}_{J}(\bar{\tau})~{}.$ (2.16)
These states $\bra{\tau,\bar{\tau}}$ | now correspond to gapless boundary conditions that define CFTs. We can also deform the theory away from the CFT fixed point. If the deformation preserves the symmetry, the bulk will be preserved, and it defines another physical boundary condition, leading to 2D partition functions that depend on the dimensionful couplings $\lambda$
$Z^{\lambda}(\tau,\bar{\tau})=\sum_{I,J}Z_{IJ}f_{I}(\lambda,\tau)\bar{f}_{J}(\lambda,\bar{\tau})=\bra{\tau,\bar{\tau},\lambda}\Psi\rangle$ (2.17)
where $f_{I}(\lambda,\tau)$ is a deformed character. For instance, the $T\bar{T}$ deformation is a notable example of a symmetry-preserving deformation [23]. Importantly, this deformation does not modify the algebraic or topological data of the 2D CFT encoded in the 3D quantum state |$\ket{\Psi}$. In the case of the $T\bar{T}$ deformation, the deformed partition functions satisfy the Callan-Symanzik equation [24, 74] and remain on the same RG trajectory. Different theories along this RG trajectory are labeled by distinct 3D states $\bra{\tau,\bar{\tau},\lambda}$ |.